a two digits number is formed using digits 3,4,5,6 and 7 without replacement .what is the probability that the number is a multiple of 3.
Answer by MathLover1(20849)

the answer is P(5, 2) = 5!/(5-2)! = 5!/3!= (5*4*3*2*1)/(3*2*1)=5*4 = 20
you can also find these two digits numbers by using a Cartesian Coordinate system if you put all numbers (3,4,5,6 and 7) on x and y axis, than find all pairs
after you find all pairs, you need to EXCLUDE pairs (3,3),(4,4),(5,5), (6,6) and (7,7)